# SurgLaVi: Large-Scale Hierarchical Dataset for Surgical Vision-Language Representation Learning

### v1

@@ 6 Conclusion
We introduce SurgLaVi, the largest and most hierarchically structured video–language dataset in the surgical domain, capturing procedural knowledge at phase, step, and task levels from narrated surgical videos. Our fully automated pipeline transforms raw surgical video collections into temporally precise, semantically rich, multi-granular annotations at unprecedented scale, setting a new benchmark for dataset design. Building on this foundation, SurgCLIP, a lightweight CLIP-style model which we design as a baseline, demonstrates that even standard architectures, when paired with high-quality hierarchical data, can achieve state-of-the-art performance across a wide spectrum of downstream tasks, from phase and step recognition to tool and action detection, in zero-shot and few-shot probing evaluation settings. These findings reveal a fundamental principle: dataset quality and structure can be more decisive than model complexity in capturing complex procedural workflows. SurgLaVi not only accelerates representation learning for surgical AI but also opens transformative opportunities for multimodal reasoning, video–text understanding, and self-supervised procedural modeling, laying the groundwork for the next generation of intelligent, context-aware surgical systems.

### v2

@@ 7 Conclusion
We introduce SurgLaVi, the largest and most hierarchically structured video–language dataset in the surgical domain, capturing procedural knowledge at coarse, mid, and fine levels from narrated surgical videos. Our fully automated pipeline transforms raw surgical video collections into temporally precise, semantically rich, multi-granular annotations at unprecedented scale, setting a new benchmark for dataset design. Building on this foundation, SurgCLIP, a lightweight CLIP-style model which we design as a baseline, demonstrates that even standard architectures, when paired with high-quality hierarchical data, can achieve state-of-the-art performance across a wide spectrum of downstream tasks, from phase and step recognition to tool and action detection, in zero-shot and few-shot probing evaluation settings. These findings reveal a fundamental principle: dataset quality and structure can be more decisive than model complexity in capturing complex procedural workflows. SurgLaVi not only accelerates representation learning for surgical AI but also opens transformative opportunities for multimodal reasoning, video–text understanding, and self-supervised procedural modeling, laying the groundwork for the next generation of intelligent, context-aware surgical systems.